# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ 3.3. Random dynamical system
In the framework of Definition 2.1, suppose that $\boldsymbol{v}(t)=\boldsymbol{y}^{\chi}(t)+\boldsymbol{z}_{\chi}(t),t\geq s,$ where $\boldsymbol{y}^{\chi}$ is the unique solution to system (3.19) with initial data $\boldsymbol{x}-\boldsymbol{z}_{\chi}(s)$ at time $s$. If the process $\{\boldsymbol{v}(t),\ t\geq s\},$ has trajectories in $\mathrm{C}([s,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([s,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$, then it is a solution to system (2.18). Vice-versa, if a process $\{\boldsymbol{v}(t),t\geq s\},$ with trajectories in $\mathrm{C}([s,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([s,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$ is a solution to system (2.18), then for any $\chi\geq 0,$ a process $\{\boldsymbol{y}^{\chi}(t),t\geq s\},$ defined by $\boldsymbol{y}^{\chi}(t)=\boldsymbol{v}(t)-\boldsymbol{z}_{\chi}(t),t\geq s,$ is a solution to (3.19) on $[s,\infty).$
4. Random attractors for stochastic third-grade fluids equations: Bounded domains
In this section, we prove the existence of unique random attractor for stochastic third-grade fluids equation 2.18 on unbounded domains. In order to prove the result of this section, we assume that external forcing term $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. Here, the RDS $\Psi$ is considered over the MDS $(\Omega,{\mathscr{F}},{\mathbb{P}},{\theta})$. The results that we have obtained in the previous sections provide a unique solution to system (2.18), which is continuous with respect to the data (particularly $\boldsymbol{x}$ and $\boldsymbol{f}$). Furthermore, if we define, for $\boldsymbol{x}\in\mathbb{H},\ \omega\in\Omega,$ and $t\geq s,$
@@ 4. Random attractors for stochastic third-grade fluids equations: Bounded domains
then the process $\{\boldsymbol{v}(t):\ t\geq s\}$ is a solution to equation (2.18)1italic-(2.18subscriptitalic-)$_{1}$ for $t>s$, for each $s\in\mathbb{R}$ and each $\boldsymbol{x}\in\mathbb{H}$.
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Suppose that $\boldsymbol{y}$ solves system (3.19) on the time interval $[a,\infty)$ with $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}(\mathbb{R};\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}(\mathbb{R};\mathbb{W}^{1,4}(\mathscr{O}))$, $\chi\geq 0$ and $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ (that is, $\boldsymbol{f}=\boldsymbol{f}_{1}+\boldsymbol{f}_{2}$ with $\boldsymbol{f}_{1}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\boldsymbol{f}_{2}\in\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$). Then, for any $t\geq\tau\geq a,$
$\|\boldsymbol{y}(t)\|^{2}_{2}$ $\leq\|\boldsymbol{y}(\tau)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}$
$\quad+C\int_{\tau}^{t}\bigg{[}\|\boldsymbol{z}(s)\|^{2}_{2}+\|\boldsymbol{z}(s)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+1\bigg{]}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-s)}\mathrm{d}s.$ (4.2)
@@
$\left|\alpha\left<\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{y}\right>\right|$ $\leq\frac{\nu(1-\varepsilon_{0})}{4}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}$ (4.5)
$\left|\beta\left<\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{z}\right>\right|$ $\leq\leq\frac{\beta\varepsilon_{0}}{12}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4},$ (4.6)
@@
and
$\big{|}\langle\boldsymbol{f},\boldsymbol{y}\rangle\big{|}$ $\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{16}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4}+1\bigg{]}.$ (4.8)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|^{2}_{2}$ $\leq-\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$\quad+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+1\bigg{]}$
$=-\nu\left(1+\frac{\varepsilon_{0}}{2}\right)\|\nabla\boldsymbol{y}(t)\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$\quad+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+1\bigg{]}$
$\leq-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\|\boldsymbol{y}(t)\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$\quad+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+1\bigg{]},$
@@
$\boldsymbol{y}(\cdot,s,\boldsymbol{x}_{m})\to\widetilde{\boldsymbol{y}}(\cdot)\ \text{ strongly in }\ \mathrm{L}^{2}(s,t;\mathbb{H}),$ (4.9)
as $m\to\infty$. Again, choosing one more subsequence (again not relabeling), we infer from (4.9) that
$\boldsymbol{y}(\tau,s,\boldsymbol{x}_{m})\to\widetilde{\boldsymbol{y}}(\tau)\text{ in }\mathbb{H},\ \text{ for a.e. }\ \tau\in[s,t],$ (4.10)
as $m\to\infty$. For $s<t$, we obtain from (4.10) that there exists $\tau\in(s,t)$ such that (4.10) holds true for this particular $\tau$. Then by Lemma 3.9, we obtain
@@
$\to\Psi\big{(}t-\tau,\theta_{\tau}\omega,\widetilde{\boldsymbol{y}}(\tau)+\boldsymbol{z}(\tau)\big{)}$
as $m\to\infty$, in $\mathbb{H}$, which completes the proof. ∎
@@
Because of [13, Theorem 2.8], it is only needed to prove that there exists a $\mathfrak{DK}$-absorbing set $\textbf{B}\in\mathfrak{DK}$ and the RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. Existence of $\mathfrak{DK}$-absorbing set $\textbf{B}\in\mathfrak{DK}$: Let $\mathrm{D}$ be a random set from the class $\mathfrak{DK}$. Let $\kappa_{\mathrm{D}}(\omega)$ be the radius of $\mathrm{D}(\omega)$, that is, $\kappa_{\mathrm{D}}(\omega):=\sup\{\|x\|_{2}:x\in\mathrm{D}(\omega)\}$ for $\omega\in\Omega$.
Let $\omega\in\Omega$ be fixed. For a given $s\leq 0$ and $\boldsymbol{x}\in\mathbb{H}$, let $\boldsymbol{y}$ be the solution of (3.19) on the time interval $[s,\infty)$ with the initial condition $\boldsymbol{y}(s)=\boldsymbol{x}-\boldsymbol{z}(s).$ Using (4.1) for $t=\xi\text{ and }\tau=s\leq-1$ with $\xi\in[-1,0]$, we obtain
@@
$\quad+C\int_{s}^{0}\bigg{\{}\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+1\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\mathrm{d}t.$ (4.11)
@@
$\quad+C\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+1\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\mathrm{d}t,$ (4.12)
@@
The above inequality implies that for $\omega\in\Omega$, $\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\in\textbf{B}(\omega)$, for all $s\leq-t_{\mathrm{D}}(\omega).$ This proves B absorbs $\mathrm{D}$. The RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. In order to establish the $\mathfrak{DK}$-asymptotically compactness of $\Psi$, we use the compact Sobolev embedding $\mathbb{V}\subset\mathbb{H}$. Let us assume that $\mathrm{D}\in\mathfrak{DK}$ and $\textbf{B}\in\mathfrak{DK}$ be such that B absorbs $\mathrm{D}$. Let us fix $\omega\in\Omega$ and take a sequence of positive numbers $\{t_{m}\}^{\infty}_{m=1}$ such that $t_{1}\leq t_{2}\leq t_{3}\leq\cdots$ and $t_{m}\to\infty$. We take an arbitrary $\mathbb{H}$-valued sequence $\{\boldsymbol{x}_{m}\}^{\infty}_{m=1}$ such that $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega),$ for all $m\in\mathbb{N}.$
Since $t_{m}\to\infty$, there exists $M_{1}=M_{1}(\tau,\hat{K})\in\mathbb{N}$ such that $t_{m}\geq t_{\mathrm{D}}(\omega)$ for all $m\geq M_{1}.$ Therefore, in view of (4.14), the sequence
@@
5. Random attractors for stochastic third-grade fluids equations: Unbounded domains
In this section, we prove the existence of unique random attractor for stochastic third-grade fluids equation 2.18 on unbounded domains. In order to prove the result of this section, we assume that external forcing term $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. The following two lemmas help us to prove the existence of random $\mathfrak{DK}$-absorbing set.

### v2

@@ 3.3. Random dynamical system
In the framework of Definition 2.1, suppose that $\boldsymbol{v}(t)=\boldsymbol{y}^{\chi}(t)+\boldsymbol{z}_{\chi}(t),t\geq s,$ where $\boldsymbol{y}^{\chi}$ is the unique solution to system (3.19) with initial data $\boldsymbol{x}-\boldsymbol{z}_{\chi}(s)$ at time $s$. If the process $\{\boldsymbol{v}(t),\ t\geq s\},$ has trajectories in $\mathrm{C}([s,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([s,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$, then it is a solution to system (2.30). Vice-versa, if a process $\{\boldsymbol{v}(t),t\geq s\},$ with trajectories in $\mathrm{C}([s,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([s,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$ is a solution to system (2.30), then for any $\chi\geq 0,$ a process $\{\boldsymbol{y}^{\chi}(t),t\geq s\},$ defined by $\boldsymbol{y}^{\chi}(t)=\boldsymbol{v}(t)-\boldsymbol{z}_{\chi}(t),t\geq s,$ is a solution to (3.19) on $s,\infty).$
4. Random attractors for a subclass of stochastic third-grade fluids equations: Bounded domains
In this section, we prove the existence of unique random attractor for system 2.30 on bounded domains. In order to prove the result of this section, we assume that external forcing term $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$. Here, the RDS $\Psi$ is considered over the MDS $(\Omega,{\mathscr{F}},{\mathbb{P}},{\theta})$. The results that we have obtained in the previous sections provide a unique solution to system (2.30), which is continuous with respect to the data (particularly $\boldsymbol{x}$ and $\boldsymbol{f}$). Furthermore, if we define, for $\boldsymbol{x}\in\mathbb{H},\ \omega\in\Omega,$ and $t\geq s,$
@@ 4. Random attractors for a subclass of stochastic third-grade fluids equations: Bounded domains
[…]
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Suppose that $\boldsymbol{y}$ solves system (3.19) on the time interval $a,\infty)$ with $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}(\mathbb{R};\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}(\mathbb{R};\mathbb{W}^{1,4}(\mathscr{O}))$, $\chi\geq 0$ and $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$. Then, for any $t\geq\tau\geq a,$
$\|\boldsymbol{y}(t)\|^{2}_{2}$ $\leq\|\boldsymbol{y}(\tau)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}+C\int_{\tau}^{t}\bigg[\|\boldsymbol{z}(s)\|^{2}_{2}+\|\boldsymbol{z}(s)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}\bigg]e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-s)}\mathrm{d}s.$ (4.2)
@@
[…]
@@
$\big|\langle\boldsymbol{f},\boldsymbol{y}\rangle\big|$ $\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+C\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}.$ (4.8)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|^{2}_{2}$ $\leq-\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}+C\bigg[\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg]$
$=-\nu\left(1+\frac{\varepsilon_{0}}{2}\right)\|\nabla\boldsymbol{y}(t)\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}+C\bigg[\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg]$
$\leq-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\|\boldsymbol{y}(t)\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}+C\bigg[\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg],$
@@
$\boldsymbol{y}(\cdot,s,\boldsymbol{x}_{m})\rightarrow\widetilde{\boldsymbol{y}}(\cdot)\ \text{ strongly in }\ \mathrm{L}^{2}(s,t;\mathbb{H}),$ (4.9)
as $m\rightarrow\infty$. Again, choosing one more subsequence (again not relabeling), we infer from (4.9) that
$\boldsymbol{y}(\tau,s,\boldsymbol{x}_{m})\rightarrow\widetilde{\boldsymbol{y}}(\tau)\text{ in }\mathbb{H},\ \text{ for a.e. }\ \tau\in[s,t],$ (4.10)
as $m\rightarrow\infty$. For $s<t$, we obtain from (4.10) that there exists $\tau\in(s,t)$ such that (4.10) holds true for this particular $\tau$. Then by Lemma 3.9, we obtain
@@
$\rightarrow\Psi\big(t-\tau,\theta_{\tau}\omega,\widetilde{\boldsymbol{y}}(\tau)+\boldsymbol{z}(\tau)\big)$
as $m\rightarrow\infty$, in $\mathbb{H}$, which completes the proof. ∎
@@
Because of [10, Theorem 2.8], it is only needed to prove that there exists a $\mathfrak{DK}$-absorbing set $\textbf{B}\in\mathfrak{DK}$ and the RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact.
Existence of $\mathfrak{DK}$-absorbing set $\textbf{B}\in\mathfrak{DK}$: Let $\mathrm{D}$ be a random set from the class $\mathfrak{DK}$. Let $\kappa_{\mathrm{D}}(\omega)$ be the radius of $\mathrm{D}(\omega)$, that is, $\kappa_{\mathrm{D}}(\omega):=\sup\{\|x\|_{2}:x\in\mathrm{D}(\omega)\}$ for $\omega\in\Omega$.
Let $\omega\in\Omega$ be fixed. For a given $s\leq 0$ and $\boldsymbol{x}\in\mathbb{H}$, let $\boldsymbol{y}$ be the solution of (3.19) on the time interval $s,\infty)$ with the initial condition $\boldsymbol{y}(s)=\boldsymbol{x}-\boldsymbol{z}(s).$ Using (4.2) for $t=\xi\text{ and }\tau=s\leq-1$ with $\xi\in[-1,0]$, we obtain
@@
$\quad+C\int_{s}^{0}\bigg\{\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}\bigg\}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\mathrm{d}t.$ (4.11)
@@
$\quad+C\int_{-\infty}^{0}\bigg\{\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{\mathbb{H}^{-1}}\bigg\}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\mathrm{d}t,$ (4.12)
@@
The above inequality implies that for $\omega\in\Omega$, $\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\in\textbf{B}(\omega)$, for all $s\leq-t_{\mathrm{D}}(\omega).$ This proves B absorbs $\mathrm{D}$.
The RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. In order to establish the $\mathfrak{DK}$-asymptotically compactness of $\Psi$, we use the compact Sobolev embedding $\mathbb{V}\subset\mathbb{H}$. Let us assume that $\mathrm{D}\in\mathfrak{DK}$ and $\textbf{B}\in\mathfrak{DK}$ be such that B absorbs $\mathrm{D}$. Let us fix $\omega\in\Omega$ and take a sequence of positive numbers $\{t_{m}\}^{\infty}_{m=1}$ such that $t_{1}\leq t_{2}\leq t_{3}\leq\cdots$ and $t_{m}\rightarrow\infty$. We take an arbitrary $\mathbb{H}$-valued sequence $\{\boldsymbol{x}_{m}\}^{\infty}_{m=1}$ such that $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega),$ for all $m\in\mathbb{N}.$
[…]
@@
5. Random attractors for a subclass of stochastic third-grade fluids equations: Unbounded domains
In this section, we prove the existence of unique random attractor for system 2.30 on unbounded domains. In order to prove the result of this section, we assume that external forcing term $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. The following two lemmas help us to prove the existence of random $\mathfrak{DK}$-absorbing set.